OpenStudy (anonymous):

solve the system 5x -5z=10 7x - 2y + 4z =35 3x +2y =-85 2) 3x + 2y - z = -7 y + 5z =-29 2x-2y = 2

OpenStudy (anonymous):

http://www.twiddla.com/865824

OpenStudy (anonymous):

question 2)reduce the first equation in two a form with 2 variable and solve the resulting equations simultaneously 2x-2y=2 2x=2+2y x=1+y substitute x in 1,you'll have 3(1+Y)+2y-z=-7 3+3y+2y-Z=-7 5y-z=-10 now solve 5y-z=-10 and y +5z=-29 you'll get y=-79/26 and x=-53/26 substitute y in y+5z=-29 to get the value of z -79/26+5z=-29 z=-135/26 try question 1

OpenStudy (anonymous):

5x -5z=10 _____1 7x - 2y + 4z =35___2 3x +2y =-85____3 3+2= 7x-2y+4z=35 3x+2y=-85 10x+4z=-50_____4 2(1) 10x-10z=20 4-1 10x+4z=-50 -10x+10z=-20 14z=-70 z=-5 now subsitute this in equation 1 5x-4(5)=10 x=6 now kindly do the same for y urself